OpenStudy (anonymous):

Solve for x: |x| − 8 = −5

OpenStudy (anonymous):

answer is 3 you add the 8 to 5 which gives you 3

OpenStudy (anonymous):

for positive charge we can write x-8=-5 =>x=-5+8 =>x=3 for negative charge we can write -(x-8)=-5 =>-x+8=-5 =>-x=-5-8 =>-x=-13 =>x=13